# Data Processing for the OpenGPT-X Model Family

### v2

@@ Abstract
This paper presents a comprehensive overview of the data preparation pipeline developed for the OpenGPT-X project, a large-scale initiative aimed at creating open and high-performance multilingual large language models (LLMs). The project goal is to deliver models that cover all major European languages, with a particular focus on real-world applications within the European Union. We explain all data processing steps, starting with the data selection and requirement definition to the preparation of the final datasets for model training. We distinguish between curated data and web data, as each of these categories is handled by distinct pipelines, with curated data undergoing minimal filtering and web data requiring extensive filtering and deduplication. This distinction guided the development of specialized algorithmic solutions for both pipelines. In addition to describing the processing methodologies, we provide an in-depth analysis of the datasets, increasing transparency and alignment with European data regulations. Finally, we share key insights and challenges faced during the project, offering recommendations for future endeavors in large-scale multilingual data preparation for LLMs.
@@ 1 Introduction
In recent years, there has been an exponential increase in the development and deployment of large language models (LLMs) [3, 42]. The most prominent and accurate models are often the products of major technology companies and are available either through APIs (e.g. OpenAI’s GPT models) or with non-open licenses (e.g. Meta’s LLaMA family [^1] Meta restricts commercial use based on monthly active users and restricts the use of LLaMa model to improve other LLMs.). Moreover, these models are often US-centric, predominantly fluent in American English, and imbued with the values inherent to that cultural context. Conversely, open and unrestricted models that focus on a variety of languages and cultures often emerge from academic settings. However, these models generally lack the accuracy and power of their more robust commercial counterparts. This disparity is primarily due to the differences in the resources available for training, such as the diversity and quality of data sources. The training data is crucial in shaping the behavior and values these models learn. The OpenGPT-X [^2] https://opengpt-x.de/en/ project aims to bridge this gap by merging the best of both worlds: delivering an open, unrestricted, high-performance model trained on a diverse dataset that covers all major European languages. OpenGPT-X targets real-world use cases, e.g. from media and the automotive sector within the European Union.
In this paper, we present a comprehensive overview of the data processing pipelines developed for the OpenGPT-X project. Our focus is solely on the collection, preparation, and curation of the data, detailing the steps for data selection, conversion, normalization, quality filtering, and deduplication. This processed data was then made available for the subsequent training of the OpenGPT-X models, such as those released in [76], which utilized a subset of the complete dataset for their development. Our main contribution are:
@@ 2.2 Multilingual datasets
OSCAR 22.01 [2] contains about 8TB of text in 153 languages. It is produced with the Ungoliant pipeline [1] from the CommonCrawl WET (Web Extracted Text) files. The filtration starts with language detection, which utilizes the fastText model to identify language of each line in a document. If the confidence score is less then 0.8, then the line is assigned with to the unknown language class. The proportion of each language in a document is calculated as a percentage of bytes in the document assigned with that language. A document is classified as multilingual if it contains at least 5 lines, not more than 5 languages, the proportion of each language is at least ${1}/{(m+1)}$ (where $m$ is the number of languages in the document) and the proportion of the unknown language is not larger than ${1}/{(m+1)}$. Otherwise, the document is classified as monolingual and is assigned with the language having the highest weighted confidence score. The weighted confidence of a language is calculated as a sum of products of byte size and language confidence of each line classified to that language divided by the total number of bytes in the document. A monolingual document is passed to further processing steps if the weighted confidence of its language is at least 0.6. Language identification is followed by the filtering of documents based on the line lengths, proportions of characters of certain Unicode classes, and the URL based UT1 blocklist. [^12] https://dsi.ut-capitole.fr/blacklists/index_en.php Deduplication is not applied to non-English data, and only line-wise deduplication is applied to English data.
@@ 3 Data selection
Given the aforementioned properties, our approach to data selection involves ensuring a large volume [^16] The amount of training data is typically measured by the number of tokens or words. However, the choice of a tokenizer affects the number of generated tokens, making absolute numbers for different models not directly comparable. Therefore, we use the wc command line utility to estimate the number of words. of diverse and high-quality processed data for LLMs training.
@@ 3.1.1 Curated data
Curated datasets are essential for providing high-quality, diverse training data for our language models. These datasets have typically undergone a quality review process, either prior to their release or through ongoing public review, and are publicly available (see Appendix 8.3 for a list of our curated datasets). This includes data such as Wikipedia, collections of scientific articles, books, and source code.
@@ 3.1.2 CommonCrawl data
Unlike curated data, web data is available in a large amount. For this reason, we prioritize processing web data only based on recency and language. Specifically, we use data from CommonCrawl (CC), an organization that maintains a free, open-source repository of web data accessible to anyone since they provide the largest amount of data. They perform monthly crawls of internet webpages, compiling them into data dumps labeled in the format YYYY-WW (e.g., 2024-22, see Section 5.2 for a list of selected dumps). CommonCrawl dumps can be downloaded from URLs [^18] Following the schema https://data.CommonCrawl.org/crawl-data/CC-MAIN-<YEAR-MONTH>/index.html. in two main formats: WARC and WET. WARC (Web ARChive format) files contain raw data from the crawl, including the full page HTML and request metadata. WET (WARC Encapsulated Text) files provide a text-only version of the websites, excluding the web metadata. In our pipeline, we use the WET files exclusively as they provide the extracted textual content of web pages in a simplified format. This eliminates the need for extensive parsing of raw HTML data found in WARC files, allowing us to focus directly on text analysis. For instance, the CC dump “2024-22” includes data crawled from 2.5 billion web pages, totaling 18.35 TiB of compressed HTML data. This particular dump consists of 90,000 WET files, which are grouped into segments for processing. This large-scale web data significantly expands the scope of our collection, providing a robust foundation for pretraining large language models.
@@ 4.1 Pipeline for curated data
Figure 1: Data pipeline for curated datasets: The pipeline involves format conversion, content normalization, followed by language detection, content normalization and filtering. The output is high-quality training data in JSONL format.
@@ 4.2 Pipeline for web data
Figure 2: Overview of the Web Data Processing Pipeline: The pipeline processes WET files with Ungoliant by normalizing, computing quality warnings and harmful-perplexity, detecting language, and prefiltering. Afterwards, data is filtered based on quality and harmful-perplexity before deduplication, producing high-quality training data in JSONL format.
Processing web data from large-scale sources like CommonCrawl presents significant challenges due to its unstructured and noisy nature. To transform this raw web data into a suitable format for training LLMs, we utilized a specialized pipeline (see Figure 2). This section provides a detailed breakdown of each processing stage, including normalization, filtering, and deduplication, using the Ungoliant pipeline.
@@ 4.2.3 Filtering
    • Iterate through the first 20% of a document [^24] The percentage is calculated on a line-base, i.e. when the character ’\$\backslash n$’ is found..
@@ 4.2.4 Document deduplication
Building on the quality filtering performed in the previous stage, the next step after running the Ungoliant pipeline is document deduplication. This process refines the dataset by removing both exact and near duplicate documents, thus reducing the overall size of the data and improving its quality for LLM training. After extensive testing [64], we found that the MinHash/LSH algorithm offers the best balance of precision, recall, and computational efficiency for deduplicating large volumes of web data. The deduplication process was executed on a per-dump and per-language basis (“local” deduplication), which is both convenient and backed by newer research [81] that confirms that local deduplication is more favorable than global (across dump and/or language) approaches for preserving data diversity while minimizing redundancy.
@@ 4.3 Technical Infrastructure
These resources enabled us to handle the extensive data processing requirements necessary for pretraining our large language models.
@@ 5 Data Analysis
It is important to note how LLM training often uses sampling techniques to optimize data distribution, considering factors like language balance and the fertility score [4]. In this paper, we focus solely on the raw data collection, filtering, and integration process, without adjusting distributions for training purposes. For further information on a model trained on part of the data we process, refer to [76].
@@ 5.1 Curated Data
Figure 3: Distribution of curated dataset licenses. Most datasets have permissive licenses (CC0, CC BY, MIT, Apache), while restrictive licenses are primarily represented by datasets sourced from multiple origins (Various).
@@ 5.1.2 Languages
(b) Word distribution excluding EN, EU24, and XX.
@@
One would expect the word distribution to follow the number of datasets available; however, correlation analysis reveals this is only partially true. The Spearman correlation is moderate (0.562, p = 0.0028), while the Pearson correlation indicates a stronger linear relationship (0.855, $P<0.001$). By using linear regression, we observe notable deviations in certain languages. For instance, German (DE) exhibits a significant negative residual, with the actual word count being 5.1 billion fewer than expected, despite having a high number of diverse datasets. This suggests that while German datasets are numerous, they tend to be smaller in size or less word-dense compared to other languages. On the other hand, English (EN) has a positive residual, showing 3.88 billion more words than predicted, which can be attributed to the oversampling of English datasets in the corpus. In contrast, French (FR) underperforms relative to its dataset count, with a negative residual of 1.54 billion, indicating that the French data is underrepresented in terms of word volume.
@@ 5.1.4 Sizes
Figure 5: Log-scaled comparison of word counts and document counts across datasets, sorted by word size. Some datasets exhibit a notably high word count despite having relatively few documents, indicating longer average document sizes.
@@ 5.1.5 Filtering
Across the curated datasets, the average filtering percentage is 5.33% (±$\pm 5.98\%$). Based on these values, filtering can be split in three categories:
  • Low Filtering ($<1\%$): Datasets that underwent minimal filtering, typically because they contained well-structured, clean data to begin with.
  • Medium Filtering (1% to 11.31%): The majority of datasets fall into this category, with moderate filtering applied to remove noise or short documents without heavily impacting the total word count.
  • High Filtering (>11.31%): Datasets with a high filtering percentage, indicating substantial noise or irrelevant content.
@@ 5.1.6 Summary
The curated dataset presents a comprehensive and diverse collection of 75 datasets spanning 25 languages, various domains, and multiple formats, offering a robust foundation for a wide range of research tasks. Despite some imbalances in distribution, particularly with large datasets dominating the total word count, the variety within the dataset makes it a valuable asset for training LLMs. The high presence of technical and legal data, alongside source code, reflects the strengths of this dataset in supporting tasks in these fields.
While filtering in general is necessary to ensure data quality, its impact on the curated datasets is relatively moderate, with only a few datasets requiring significant cleaning. The weak correlations found between filtering percentages and dataset characteristics suggest that filtering is driven more by the specificities of the dataset than by its size, format, or domain. In preparing data for training LLMs, we found that a strategic approach to selecting subsets is key. Leveraging the wide domain and language diversity can help ensure more balanced outcomes.
@@ 6 Insights and Observations
During the process of preparing the data for LLM training, several challenges emerged that shaped the development of the data pipeline. These challenges are important for understanding both the complexity of large-scale multilingual dataset processing and the technical, organizational, and legal obstacles one can encounter. By presenting these observations, we provide key findings and recommendations to guide future projects and to navigate similar challenges.
@@ 7 Conclusion and Outlook
In this paper, we presented the entire process of data preparation for the OpenGPT-X model family, covering every step from the initial selection of data to delivering the finalized datasets for model training. We outlined the requirements that guided our data selection, which conceptually split our sources into two categories: curated data and web data. This distinction shaped our approach to the implementation of our data pipelines: one for curated data, which required minimal filtering, and another for web data, which focused heavily on filtering and deduplication. In addition to providing a thorough description of our pipeline and data preparation methods, we included an in-depth analysis of the resulting datasets. This analysis ensures transparency, aligning with European data recommendations and best practices in scientific research.
Finally, to contribute to the broader research field and promote openness, we highlighted the challenges faced during this project and the lessons learned. By sharing these insights, we aim to provide valuable guidance for future projects that undertake large-scale data preparation for multilingual language models. We utilized a subset of the preprocessed datasets detailed in this paper to pretrain large language models for the OpenGPTX project.
@@ 8.2 Metadata Content Example
    "text": "Welcome to the world’s foremost open contentOrganic
@@ 8.3.1 Filtering
Finally, domain showed no significant correlation with the percentage of data filtered ($r=-0.004$, p = 0.97).
@@
8.3.2 List of Curated Datasets
@@
The complete list of curated dataset is presented in Table LABEL:tab:curated_data_list (next page).
@@
Table 4: This table provides an overview of various multilingual datasets utilized in OpenGPT-X project. Each entry begins with the Corpus ID of the dataset and a link to its project page. The Language/s column specifies the languages included within each dataset (see Section 5.1). The Format of the datasets is also noted, indicating the file type, such as TXT, JSON, or XML. The License column outlines the legal terms governing the use of each dataset, for uncommon licenses a link is provided. The Domain column reflects the specific field or subject area that the dataset pertains to, such as Law, Math, or Medical. The # Docs column presents the total number of documents contained in each dataset, while the # Words column conveys the total word count in thousand. Finally, the % Filtered (Filt.) column indicates the percentage of documents that have been filtered out during preprocessing.
Corpus ID Language/s Format License Domain #Docs #Words % Filt.
eu_admin EU24 TXT CC BY-NC 3.0 Law & Admin. 435.318 1.079 0%
ampsmath_khan [44] EN JSON / TXT MIT Math 102.985 12.081 0%
ampsmath_mathematica [44] EN JSON / TXT MIT Math 4.824.189 232.502 0%
cdrs_bt [21] DE TXT CC0 1.0 Law & Admin. 124.487 515.053 8%
ce_bag [20] DE TXT CC0 1.0 Law & Admin. 5.290 18.487 0%
ce_bfh [29] DE TXT CC0 1.0 Law & Admin. 10.310 22.635 0%
ce_bgh [25] DE TXT CC0 1.0 Law & Admin. 71.976 104.002 0%
ce_bpatg [26] DE TXT CC0 1.0 Law & Admin. 30.772 66.029 0%
ce_bverfg [23] DE TXT CC0 1.0 Law & Admin. 8.449 20.679 0%
ce_bverwg [27] DE TXT CC0 1.0 Law & Admin. 27.099 49.664 0%
cpp_bt [22] DE TXT CC0 1.0 Law & Admin. 4.087 235.502 0%
gaois_corpus GA XML / TMX CC BY 4.0 Law & Admin. 2 2.168 0%
cd_icj [28] EN, FR TXT CC0 1.0 Law & Admin. 4.565 25.979 0%
cd_pcij [24] EN TXT CC0 1.0 Law & Admin. 518 2.117 0%
dagw [17] DA TXT CC BY 4.0 Web 285.634 922.700 5%
dta DE TXT CC BY-SA 4.0 Books 5.305 197.339 0%
dcep [73] EU24 TXT EUPL Law & Admin. 1.227.996 1.117 5%
medi_notice [36] DE, FR, IT TSV CC BY SA Medical 22.219 52.068 0%
enc2021 [55] EU24 VERT CC BY-NC 4.0 Web 4.361.539 1.004 5%
estonian_reference_corpus ET XML CC BY-NC Web 17.892 227.978 0%
emea [89] EU24 TXT CC BY 4.0 Medical 17.960 152.252 4%
europarl [54] EU24 XML CC0 1.0 Law & Admin. 63.937 660.716 0%
eurlex [62] EU24 JSONL CC BY 4.0 Law & Admin. 4.463.480 13.235 8%
ecb_corpus [89] EU24 XML CC0 Finance 19 53.894 0%
fi_news FI CSV MIT News 69.413 3.650 62% [^27] Very short documents (headlines)
german_legal_cases [78] DE JSONL CC0 1.0 Law & Admin. 249.240 749.060 0%
german_political_speeches [6] DE XML CC BY-SA 4.0 Law & Admin. 6.659 11.366 0%
greek_legal_code [79] EL JSONL CC BY 4.0 Law & Admin. 40.929 28.791 1%
irish_legislation GA XML / TMX CC BY 4.0 Law & Admin. 12 29.136 0%
kleine_anfragen DE SQL ODbL 1.0 Law & Admin. 3.401.475 6.151 18% [^28] Noise from PDF conversion.
korpus_malti [71] MA JSONL CC BY-NC-SA Web 103.874 276.204 11% [^29] Noisy data and short document.
legal_mc4 EU24 JSONL CC BY 4.0 Law & Admin. 4.853.498 13.892 1%
marcell [90] BG, HR, HU, RO, SL, SK, PL CONLLUP / TXT CC0 1.0 Law & Admin. 280.629 373.283 5%
macocu [9] EU24 XML CC0 1.0 Web 79.164.225 27.930 12% [^30] Short documents.
multi_un [18] EU24 XML CC BY-NC Law & Admin. 262.391 1.096 0%
multi_eurlex [38] EU24 JSONL CC BY-NC-SA 4.0 Law & Admin. 1.038.305 1.184 1%
medical_t5 [34] DE, EN, ES, FR JSON Apache-2.0 Medical 94.269 90.741 1%
@@
ncc [60] NO JSON Various [^31] The ncc corpus includes the following licenses : NLOD 2.0, CC0 1.0, CC BY-NC 2.0, CC BY-SA 3.0 Knowledge Base 1.241.694 2.920 60% [^32] Short documents.
openlegaldata [78] DE JSONL CC0 1.0 Law & Admin. 103.870 363.962 0%
opensubtitles2018 [66] EU24 XML CC BY-NC Culture 2.257.513 1.398 22% [^33] Noisy documents.
open_web_math [80] EN Parquet ODC-By 1.0 Math 5.956.611 7.378 2%
paisa [68] IT TXT / XML CC BY-NC-SA 3.0 Web 318.730 207.382 7%
parlamint [19] EU24 XML / TEI CC BY Law & Admin. 40.995 1.095 3%
pes2o [87] EN JSONL ODC-By 1.0 Academic 38.958.175 42.172 0%
pile_v2_freelawopinions [32] EN JSONL CC BY-ND 4.0 Law & Admin. 4.410.012 10.408 0%
pile_hackernews [32] EN JSONL CC BY-NC Web 331.205 246.149 1%
pile_nih_exporter [32] EN JSONL CC0 1.0 Medical 930.048 307.055 0%
pile_openwebtext2 [32] EN JSONL MIT Books 15.752.217 10.633 1%
@@
pile_v2_philarchive [32] EN JSONL Various [^34] This Pile part is extracted from ArXiv where the author can decide between a variety of licenses. Academic 40.344 487.032 1%
@@
pile_pmc_abstracts [32] EN JSONL Various [^35] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 15.476.085 3.173 0%

### v4

@@ Abstract
This paper presents a comprehensive overview of the data preparation pipeline developed for the OpenGPT-X project, a large-scale initiative aimed at creating open and high-performance multilingual large language models (LLMs). The project goal is to deliver models that cover all major European languages, with a particular focus on real-world applications within the European Union. We explain all data processing steps, starting with the data selection and requirement definition. We distinguish between curated data and web data, as each of these categories is handled by distinct pipelines, with curated data undergoing minimal filtering and web data requiring extensive filtering and deduplication. This distinction guided the development of specialized algorithmic solutions for both pipelines. In addition to describing the processing methodologies, we provide an in-depth analysis of the datasets, increasing transparency and alignment with European data regulations. Finally, we share key insights and challenges faced during the project, offering recommendations for future endeavors in large-scale multilingual data preparation for LLMs.
@@ 1 Introduction
In recent years, there has been an exponential increase in the development and deployment of large language models (LLMs) [3, 42]. The most prominent and accurate models are often the products of major technology companies and are available either through APIs (e.g. OpenAI’s GPT models) or with non-open licenses (e.g. Meta’s LLaMA family [^1] Meta restricts commercial use based on monthly active users and restricts the use of LLaMa model to improve other LLMs.). Moreover, these models are often US-centric, predominantly fluent in American English, and imbued with the values inherent to that cultural context. Conversely, open and unrestricted models that focus on a variety of languages and cultures often emerge from academic settings. However, these models generally lack the accuracy and power of their more robust commercial counterparts. This disparity is primarily due to the differences in the resources available for training, such as the diversity and quality of data sources. The training data is crucial in shaping the behavior and values these models learn. The OpenGPT-X [^2] https://opengpt-x.de/en/ project aims to bridge this gap by merging the best of both worlds: delivering an open, unrestricted, high-performance model trained on diverse data that covers all major European languages. OpenGPT-X targets real-world use cases, e.g. from media and the automotive sector within the European Union.
In this paper, we present a comprehensive overview of the data processing pipelines developed for the OpenGPT-X project. Our focus is solely on the collection, preparation, and curation of the data, detailing the steps for data selection, conversion, normalization, quality filtering, and deduplication. The processed data was then made available for training of OpenGPT-X models [76], which utilized and sampled a subset of the data. Our main contribution are:
@@ 2.2 Multilingual datasets
OSCAR 22.01 [2] contains about 8TB of text in 153 languages. It is produced with the Ungoliant pipeline [1] from the CommonCrawl WET (Web Extracted Text) files. The filtration starts with language detection, which utilizes the fastText model to identify language of each line in a document. If the confidence score is less then 0.8, then the line is assigned with to the unknown language class. The proportion of each language in a document is calculated as a percentage of bytes in the document assigned with that language. A document is classified as multilingual if it contains at least 5 lines, not more than 5 languages, the proportion of each language is at least 1/(m+1){1}/{(m+1)} (where $m$ is the number of languages in the document) and the proportion of the unknown language is not larger than 1/(m+1){1}/{(m+1)}. Otherwise, the document is classified as monolingual and is assigned with the language having the highest weighted confidence score. The weighted confidence of a language is calculated as a sum of products of byte size and language confidence of each line classified to that language divided by the total number of bytes in the document. A monolingual document is passed to further processing steps if the weighted confidence of its language is at least 0.6. Language identification is followed by the filtering of documents based on the line lengths, proportions of characters of certain Unicode classes, and the URL based UT1 blocklist. [^12] https://dsi.ut-capitole.fr/blacklists/index_en.php Deduplication is not applied to non-English data, and only line-wise deduplication is applied to English data.
@@ 3 Data selection
Given the aforementioned properties, we suggest a method for data selection that ensures a large volume [^16] The amount of training data is typically measured by the number of tokens or words. However, the choice of a tokenizer affects the number of generated tokens, making absolute numbers for different models not directly comparable. Therefore, we use the wc command line utility to estimate the number of words. of diverse and high-quality after preprocessing.
@@ 3.1.1 Curated data
Curated datasets are essential for providing high-quality, diverse training data for LLMs. These datasets have typically undergone a quality review process, either prior to their release or through ongoing public review, and are publicly available (see Appendix 8.3 for a list of curated datasets we compiled). The list includes data such as Wikipedia, collections of scientific articles, books, and source code.
@@ 3.1.2 CommonCrawl data
Unlike curated data, web data is available in a large amount. For this reason, we prioritize processing web data only based on recency and language. Specifically, we use data from CommonCrawl (CC), an organization that maintains a free, open-source repository of web data accessible to anyone since they provide the largest amount of data. They perform monthly crawls of internet webpages, compiling them into data dumps labeled in the format YYYY-WW (e.g., 2024-22, see Section 5.2 for a list of selected dumps). CommonCrawl dumps can be downloaded from URLs [^18] Following the schema https://data.CommonCrawl.org/crawl-data/CC-MAIN-<YEAR-MONTH>/index.html. in two main formats: WARC and WET. WARC (Web ARChive format) files contain raw data from the crawl, including the full page HTML and request metadata. WET (WARC Encapsulated Text) files provide a text-only version of the websites, excluding the web metadata. In our pipeline, we use the WET files exclusively as they provide the extracted textual content of web pages in a simplified format. This eliminates the need for extensive parsing of raw HTML data found in WARC files, allowing us to focus directly on text analysis. For instance, the CC dump “2024-22” includes data crawled from 2.5 billion web pages, totaling 18.35 TiB of compressed HTML data. This particular dump consists of 90,000 WET files, which are grouped into segments for processing.
@@ 4.1 Pipeline for curated data
Figure 1: Data pipeline for curated datasets: The pipeline involves format conversion, content normalization, followed by language detection, content normalization and filtering. The output is high-quality preprocessed data in JSONL format.
@@ 4.2 Pipeline for web data
Figure 2: Overview of the Web Data Processing Pipeline: The pipeline processes WET files with Ungoliant by normalizing, computing quality warnings and harmful-perplexity, detecting language, and prefiltering. Afterwards, data is filtered based on quality and harmful-perplexity before deduplication, producing high-quality preprocessed data in JSONL format.
Processing web data from large-scale sources like CommonCrawl presents significant challenges due to its unstructured and noisy nature. To transform the raw web data into a suitable format for training LLMs, we utilized a specialized pipeline (see Figure 2). This section provides a detailed breakdown of each processing stage, including normalization, filtering, and deduplication, using the Ungoliant pipeline.
@@ 4.2.3 Filtering
[…]
@@ 4.2.4 Document deduplication
Building on the quality filtering performed in the previous stage, the next step after running the Ungoliant pipeline is document deduplication. Deduplication refines the dataset by removing both exact and near duplicate documents, thus reducing the overall size of the data. After extensive testing [64], we found that the MinHash/LSH algorithm offers the best balance of precision, recall, and computational efficiency for deduplicating large volumes of web data. The deduplication process was executed on a per-dump and per-language basis (“local” deduplication), which is both convenient and backed by newer research [81] that confirms that local deduplication is more favorable than global (across dump and/or language) approaches for preserving data diversity while minimizing redundancy.
@@ 4.3 Technical Infrastructure
These compute resources enabled us to handle the extensive data processing requirements necessary for pretraining our large language models.
@@ 5 Data Analysis
It is important to note how LLM training often uses sampling techniques to optimize data distribution, considering factors like language balance and the fertility score [4]. In this paper we focus on techniques to process raw data into high-quality data for subsequent steps in LLM training, e.g. sampling.
@@ 5.1 Curated Data
Figure 3: Distribution of curated dataset licenses. Most datasets have permissive licenses (CC0, $C$ BY, MIT, Apache), while restrictive licenses are primarily represented by datasets sourced from multiple origins (Various).
@@ 5.1.2 Languages
(b) Word distribution excluding EN, EU24, and $X$.
@@
One would expect the word distribution to follow the number of datasets available; however, correlation analysis reveals this is only partially true. The Pearson correlation indicates a stronger linear relationship (0.855, $P<0.001$). By using linear regression, we observe notable deviations in certain languages. For instance, German (DE) exhibits a significant negative residual, with the actual word count being 5.1 billion fewer than expected, despite having a high number of diverse datasets. This suggests that while German datasets are numerous, they tend to be smaller in size or less word-dense compared to other languages. On the other hand, English (EN) has a positive residual, showing 3.88 billion more words than predicted, which can be attributed to the oversampling of English datasets in the corpus. In contrast, French (FR) underperforms relative to its dataset count, with a negative residual of 1.54 billion, indicating that the French data is underrepresented in terms of word volume.
@@ 5.1.4 Sizes
Figure 5: Log-scaled comparison of word counts and document counts across datasets, sorted by word size. Some datasets exhibit a notably high word count despite having relatively few documents, indicating longer than average document sizes.
@@ 5.1.5 Filtering
Across the curated datasets, the average filtering percentage is 5.33% ($\pm 5.98\%$). Based on these values, filtering can be split in three categories:
  • Low Filtering ($1\%$): Datasets that underwent minimal filtering, typically because they contained well-structured, clean data to begin with.
  • Medium Filtering (1% to 15%): The majority of datasets fall into this category, with moderate filtering applied to remove noise or short documents without heavily impacting the total word count.
  • High Filtering (>15%): Datasets with a high filtering percentage, indicating substantial noise or irrelevant content.
@@ 5.1.6 Summary
The curated data presents a comprehensive and diverse collection of 75 datasets spanning 25 languages, various domains, and multiple formats, offering a robust foundation for a wide range of research tasks. The variety within the data makes it a valuable asset even though there are imbalances in the distributions originating from large datasets dominating the total word count. These imbalances can be remedied by subsequent sampling techniques to balance the data. The prevalence of technical and legal data, alongside source code, reflects the strengths of the preprocessed data in supporting tasks in these fields.
While filtering in general is necessary to ensure data quality, its impact on the curated datasets is relatively moderate, with only a few datasets requiring significant cleaning. The weak correlations found between filtering percentages and dataset characteristics suggest that filtering is driven more by the specificity of the dataset than by its size, format, or domain. We found that a strategic approach to data selection is a key challenge. Leveraging the wide domain and language diversity can help ensure more balanced outcomes.
@@ 6 Insights and Observations
During the process of data preprocessing, several challenges emerged that shaped the development of the data pipeline. These challenges are important for understanding both the complexity of large-scale multilingual dataset processing and the technical, organizational, and legal obstacles one can encounter. By presenting these observations, we provide key findings and recommendations to guide future projects and to navigate similar challenges.
@@ 7 Conclusion and Outlook
In this paper, we presented the entire process of data preparation for the OpenGPT-X project. We outlined the requirements that guided our data selection, which conceptually split our sources into two categories: curated data and web data. This distinction shaped our approach to the implementation of our data pipelines: one for curated data, which required minimal filtering, and another for web data, which focused heavily on filtering and deduplication. In addition to providing a thorough description of our pipeline and data preparation methods, we included an in-depth analysis of the resulting datasets. This analysis ensures transparency, aligning with European data recommendations and best practices in scientific research.
Finally, to contribute to the broader research field and promote openness, we highlighted the challenges faced during this project and the lessons learned. By sharing these insights, we aim to provide valuable guidance for future projects that undertake large-scale data preparation for multilingual language models.
@@ 8.2 Metadata Content Example
    "text": "Welcome to the world’s foremost open content
@@ 8.3.1 Filtering
Finally, domain showed no significant correlation with the percentage of data filtered (r=0.004r=-0.004, p = 0.97).
@@
List of Curated Datasets
@@
The complete list of curated datasets is presented in Table LABEL:tab:curated_data_list (next page).
@@
Table 4: This table provides an overview of various multilingual datasets utilized in OpenGPT-X project. Each entry begins with the Corpus ID of the dataset and a link to its project page. The Language(s) column specifies the languages included within each dataset (see Section 5.1). The Format of the datasets is also noted, indicating the file type, such as TXT, JSON, or XML. The License column outlines the legal terms governing the use of each dataset, for uncommon licenses a link is provided. The Domain column reflects the specific field or subject area that the dataset pertains to, such as Law, Math, or Medical. The #Docs column presents the total number of documents contained in each dataset, while the #Words column conveys the total word count in thousand.
Corpus ID Language(s) Format License Domain #Docs #Words
eu_admin EU24 TXT CC BY-NC 3.0 Law & Admin. 435.318 1.079
ampsmath_khan [44] EN JSON / TXT MIT Math 102.985 12.081
ampsmath_mathematica [44] EN JSON / TXT MIT Math 4.824.189 232.502
cdrs_bt [21] DE TXT CC0 1.0 Law & Admin. 124.487 515.053
ce_bag [20] DE TXT CC0 1.0 Law & Admin. 5.290 18.487
ce_bfh [29] DE TXT CC0 1.0 Law & Admin. 10.310 22.635
ce_bgh [25] DE TXT CC0 1.0 Law & Admin. 71.976 104.002
ce_bpatg [26] DE TXT CC0 1.0 Law & Admin. 30.772 66.029
ce_bverfg [23] DE TXT CC0 1.0 Law & Admin. 8.449 20.679
ce_bverwg [27] DE TXT CC0 1.0 Law & Admin. 27.099 49.664
cpp_bt [22] DE TXT CC0 1.0 Law & Admin. 4.087 235.502
gaois_corpus GA XML / TMX CC BY 4.0 Law & Admin. 2 2.168
cd_icj [28] EN, FR TXT CC0 1.0 Law & Admin. 4.565 25.979
cd_pcij [24] EN TXT CC0 1.0 Law & Admin. 518 2.117
dagw [17] DA TXT CC BY 4.0 Web 285.634 922.700
dta DE TXT CC BY-SA 4.0 Books 5.305 197.339
dcep [73] EU24 TXT EUPL Law & Admin. 1.227.996 1.117
medi_notice [36] DE, FR, IT TSV CC BY SA Medical 22.219 52.068
enc2021 [55] EU24 VERT CC BY-NC 4.0 Web 4.361.539 1.004
estonian_reference_corpus ET XML CC BY-NC Web 17.892 227.978
emea [89] EU24 TXT CC BY 4.0 Medical 17.960 152.252
europarl [54] EU24 XML CC0 1.0 Law & Admin. 63.937 660.716
eurlex [62] EU24 JSONL CC BY 4.0 Law & Admin. 4.463.480 13.235
ecb_corpus [89] EU24 XML CC0 Finance 19 53.894
fi_news FI CSV MIT News 69.413 3.650
german_legal_cases [78] DE JSONL CC0 1.0 Law & Admin. 249.240 749.060
german_political_speeches [6] DE XML CC BY-SA 4.0 Law & Admin. 6.659 11.366
greek_legal_code [79] EL JSONL CC BY 4.0 Law & Admin. 40.929 28.791
irish_legislation GA XML / TMX CC BY 4.0 Law & Admin. 12 29.136
kleine_anfragen DE SQL ODbL 1.0 Law & Admin. 3.401.475 6.151
korpus_malti [71] MA JSONL CC BY-NC-SA Web 103.874 276.204
legal_mc4 EU24 JSONL CC BY 4.0 Law & Admin. 4.853.498 13.892
marcell [90] BG, HR, HU, RO, SL, SK, PL CONLLUP / TXT CC0 1.0 Law & Admin. 280.629 373.283
macocu [9] EU24 XML CC0 1.0 Web 79.164.225 27.930
multi_un [18] EU24 XML CC BY-NC Law & Admin. 262.391 1.096
multi_eurlex [38] EU24 JSONL CC BY-NC-SA 4.0 Law & Admin. 1.038.305 1.184
medical_t5 [34] DE, EN, ES, FR JSON Apache-2.0 Medical 94.269 90.741
@@
ncc [60] NO JSON Various [^27] The ncc corpus includes the following licenses : NLOD 2.0, CC0 1.0, CC BY-NC 2.0, CC BY-SA 3.0 Knowledge Base 1.241.694 2.920
openlegaldata [78] DE JSONL CC0 1.0 Law & Admin. 103.870 363.962
opensubtitles2018 [66] EU24 XML CC BY-NC Culture 2.257.513 1.398
open_web_math [80] EN Parquet ODC-By 1.0 Math 5.956.611 7.378
paisa [68] IT TXT / XML CC BY-NC-SA 3.0 Web 318.730 207.382
parlamint [19] EU24 XML / TEI CC BY Law & Admin. 40.995 1.095
pes2o [87] EN JSONL ODC-By 1.0 Academic 38.958.175 42.172
pile_v2_freelawopinions [32] EN JSONL CC BY-ND 4.0 Law & Admin. 4.410.012 10.408
pile_hackernews [32] EN JSONL CC BY-NC Web 331.205 246.149
pile_nih_exporter [32] EN JSONL CC0 1.0 Medical 930.048 307.055
pile_openwebtext2 [32] EN JSONL MIT Books 15.752.217 10.633
@@
pile_v2_philarchive [32] EN JSONL Various [^28] This Pile part is extracted from ArXiv where the author can decide between a variety of licenses. Academic 40.344 487.032
@@
pile_pmc_abstracts [32] EN JSONL Various [^29] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 15.476.085 3.173